Common fixed point theorems with applications in dynamic programming H. K. Pathak Bhilai Nagar, India and Brian Fisher, Leicester, England

Keywords: Common fixd point, coincidence point, dynamic programming AMS Subject classification: 54H25, 47H10 Abstract: Some common fixed point theorems for two pairs of mappings are proved and some applications are given in dynamic programming.